OpenStudy (anonymous):

Hello, I am like NOT understanding linear, proportional, inversely proportional, non linear equations and graphs. SOMEONE HELP!

OpenStudy (anonymous):

what is it that you don't understand?

OpenStudy (anonymous):

um,,.... everything. :( how do u no which is which?

OpenStudy (anonymous):

i no linear must have an exponent in the equation, and that linear must go through (0,0) but the rest i dont know

OpenStudy (anonymous):

a linear equation represents a straight line on a graph. any equation of the form ax+by = c is a linear equation.

OpenStudy (anonymous):

um. im in 7th grade. we learned y=mx+b

OpenStudy (anonymous):

yes, that is the same as ax+by = c, where a = m, b = -1 and c = -b

OpenStudy (anonymous):

we learned, y=slope times X plus or minus y intercept

OpenStudy (anonymous):

you are right ax+by = c is another way of representing the same equation.

OpenStudy (anonymous):

o. thx. my teacher says that exponential (non linear) uses an exponent in the equation.... how do u find it on a graph??? my TCAP is tomoro!

OpenStudy (anonymous):

so a linear equation represents a line, i.e y = mx+b is a linear equation. it is called a linear equation because as x changes, y changes linearly. a non linear equation has an exponent. say y =x^2. x has an exponent 2. which means if you plot this graph, it won't be a straight line.

OpenStudy (anonymous):

but neither would an inversely proportional line. if it goes through two lines its inversely right? otherwise its not linear

OpenStudy (anonymous):

http://www.school-for-champions.com/algebra/images/linear_equations-2_variable.gif that is a linear equation.

OpenStudy (anonymous):

a proportional equation means that y is proportional to x, i.e y =cx, where c is a constant.

OpenStudy (anonymous):

inversely proportional means as x increases, y decreases. i.e y is proportional to 1/x or y = c/x, where c is a constant.

OpenStudy (anonymous):

so it doesnt go through 0,0? OH! C or c over x. i thought that was inversely? X

OpenStudy (anonymous):

yeah in an equation y=mx+b is linear.... an exponent is non linear.... c over X is inversely..... and proportional is?????

OpenStudy (anonymous):

for an inversely proportional graph, if x was 0, y would be infinity. so no, inversely proportional graphs dont go through 0,0

OpenStudy (anonymous):

which is a parabala? cathryncausey94.blogspot.com/2010_12_01_archi...

OpenStudy (anonymous):

okay lets take y = mx+b ------> linear equation if b = 0, i.e , y = mx -----------> proportional equation. if x and/or y has exponent other than 1, non linear so if exponent of x = -1, b =0 y = mx^-1---------> y = m/x-------> inversely proportional

OpenStudy (anonymous):

is that clear?

OpenStudy (anonymous):

O... proportional is when the y intercept is 0 or not their! :) thanx.... in a functions table how would i know which was inversely and which was proportional

OpenStudy (anonymous):

think about what the word proportional means. Suppose i tell you my profits are proportional to my sales. what does that mean? the more sales I have, the more profit I make. similarly , y = cx means that y is proportional to x. the constant c defines how proportional it is.

OpenStudy (anonymous):

... im sorry. i dont speak math. :( ummm. in a functions table such as.... an input output table how would u no which was which? something about dividing and multiplying.

OpenStudy (anonymous):

its not math, it is plain english. anyway. if you see that y = cx, then it is proportional. if y = c/x, it is inversely proportional